Groundwater contains soluble Fe2+. Using standard free energies and enthalpies of formation, determine the equilibrium constant (K) at 25
Answer #1

deltaG and equilibrium constant is related as deltaG = -RT lnK

Fe2+ = Fe3+ + e +74.27

(1/4O2) + (H+) + e- = 1/2H2O; -118.59 kJ/mol,

adding these two reactions, gives deltaG = -44.27 and solving K in the above equation, we get K = 1.02

Since water is present in excess quantity, it can be eliminated from the equation and so can be the molecular oxygen as it is gas and hydrogen ion, as the pH is kept at 7.

K = [Fe3+][H2O]^1/2 / [Fe2+][O2]^1/4[H+] = [Fe3+]/[Fe2+] = 1.02